# Recursive subhomogeneity of orbit-breaking subalgebras of $\mathrm{C}^{*}$-algebras associated to minimal homeomorphisms twisted by line bundles

### v1

@@ 6.0.2. Local trivialization of $\mathscr{M}^{(n)}$
    since $\mathscr{V}^{(i)}_{x}$ and $\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ are all 1-dimensional. Using (13) we see that
    $\varsigma^{(i)}(x)=s_{m+i,i}(x)v^{(m)}_{\alpha^{i}(U)}(\alpha^{i}(x)),$ (15)
    since
    $v^{(m)}_{\alpha^{i}(U)}(\alpha^{i}(x))={h_{\alpha^{m-1}(\alpha^{i}(U))}}(\alpha^{m-1}(\alpha^{i}(x)),1)\otimes\cdots\otimes{h_{\alpha^{i}(U)}}(\alpha^{i}(x),1).$
Note that $\Gamma(\mathscr{M}^{(n)})$ is an algebra with respect to multiplication defined by pointwise composition of endomorphisms at each fibre. For $\varsigma\in\Gamma(\mathscr{M}^{(n)})$ and $x\in U\cap V$, we have
$H_{U}^{(n)}(x,(H_{U}^{(n)})^{-1}(\varsigma(x))^{*})$ $=\ H_{U}^{(n)}(x,\widetilde{\varsigma}_{U}(x)^{*})$
$=\ H_{V}^{(n)}(x,u_{VU}\widetilde{\varsigma}_{U}(x)^{*}u_{VU}^{*})$
$=\ H_{V}^{(n)}(x,(u_{VU}\widetilde{\varsigma}_{U}(x)u_{VU})^{*})$
$=\ H_{V}^{(n)}(x,\widetilde{\varsigma}_{V}(x)^{*})$
$=\ H_{V}^{(n)}(x,(H_{V}^{(n)})^{-1}(\varsigma(x))^{*}),$
hence $\varsigma^{*}$ is well-defined by $\varsigma^{*}(x):=H_{U}^{(n)}(x,(H_{U}^{(n)})^{-1}(\varsigma(x))^{*})$. The -algebra $\Gamma(\mathscr{M}^{(n)})$ is in fact a $C^{*}$-algebra in the norm
$\|\varsigma\|:=\sup_{x\in X}\|(H_{U}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}.$
Note that $\|(H_{U}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}=\|(H_{V}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}$, so the norm above does not depend on the choice of the set $U$.
From now on, we will assume that we work with a fixed minimal homeomorphism $\alpha$ on a compact metric space $X$, a fixed line bundle $\mathscr{V}=[V,p,X]$ and a fixed closed subset $Y\subseteq X$ with non-empty interior.
The purpose of this paper is to show that the associated orbit-breaking algebra $\mathcal{O}(C_{0}(Y\setminus X)\Gamma(\mathscr{V}))$ has recursive subhomogeneous algebra structure. Since the proof is quite long, we will describe the main steps here.
We will denote $\Gamma(\mathscr{V},\alpha)$ by $\mathcal{E}$ and let $\mathcal{E}_{Y}$ be the orbit breaking bimodule $C_{0}(X\setminus Y)\mathcal{E}$.
@@
For $k=1,2,\ldots,K$, we define $\mathscr{M}_{k}$ to be the restriction of the endomorphism bundle $\mathscr{M}^{(r_{k})}$ onto $\overline{Y_{k}}$, namely
@@
For convenience, we write $H_{U}^{(k)}:=H_{U}^{(r_{k})}$, $U\in\mathcal{U}$, for the rest of the paper.
In this section we show that there exists an embedding $\pi$ of the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$, where $\Gamma(\mathscr{M}_{k})$ is the $C^{*}$-algebra of continuous sections of the endomorphism bundle $\mathscr{M}_{k}$. We begin with constructing a $\mathrm{C}^{*}$-homomorphism $\pi_{k}$ from $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ for every $k=1,\dots,K$. Then we set $\pi=\pi_{1}\oplus\dots\oplus\pi_{K}$. We end this section with proving that $\pi_{1}$ is surjective while it is not true for $\pi_{k}$ when $2\leq k\leq K$ unless $Y_{k}$ is closed.
To construct a *-homomorphism $\pi_{k}$ from the Cuntz–Pimsner algebra $\mathcal{O}(\mathcal{E}_{Y})$ into the $\mathrm{C}^{*}$-algebra $\Gamma(\mathscr{M}_{k})$, we define a covariant representation $(\pi_{k},\tau_{k})$ of $\mathcal{E}_{Y}$ into $\mathrm{C}^{*}$-algebra $\Gamma(\mathscr{M}_{k})$ which gives rise to the integrated -homomorphism of $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ which we will also denote by $\pi_{k}$.
@@ 7. Embedding $\mathcal{O}(\mathcal{E}_{Y})$ into a direct sum of homogeneous algebras
$\pi_{k}(f)(x)(\mathbf{a})$ $=(f(x)a_{0},f(\alpha(x))a_{1},\ldots,f(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
$\tau_{k}(\xi)(x)(\mathbf{a}),$ $=(0,\xi(x)\otimes a_{0},\xi(\alpha(x))\otimes a_{1},\ldots,\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
[…]
Let $x\in Y_{k}\cap U$, $U\in\mathcal{U}$, and consider the basis $\{e_{U,0}(x),\dots,e_{U,r_{k}-1}(x)\}$ of $\mathscr{D}^{(r_{k})}_{x}$. If $\xi\in\mathcal{E}_{Y}$, then for each $j=0,\dots,r_{k}-2$, the scalar $\widetilde{\xi}(\alpha^{j}(x))\in\mathbb{C}$ with
$\xi(\alpha^{j}(x))=\widetilde{\xi}(\alpha^{j}(x))h_{\alpha^{j}(U)}(\alpha^{j}(x),1),$ (16)
@@
$\tau_{k}(\xi)(x)e_{U,j}(x)=\widetilde{\xi}(\alpha^{j}(x))e_{U,j+1}(x),$ (17)
@@
$U_{x}(\lambda_{0},\dots,\lambda_{r_{k}-1})=\lambda_{0}e_{U,0}(x)+\cdots+\lambda_{r_{k}-1}e_{U,r_{k}-1}(x),$
used above, we have the following matrix representation
@@
This matrix representation will be used in the subsequent discussions along with the fact that the matrix depends on the choice of the open set $U$ containing $x$, but the matrix entries differ only by multiplication by scalars of modulus one. For $f\in C(X)$, the corresponding matrix representation is clearly
@@
It is rather obvious that $\pi_{k}$ is a -homomorphism, and we may write $\pi_{k}(f)(x)$ as
@@
where $\xi(x)^{*}\otimes a_{1}:=\langle\xi(x),a_{1}\rangle_{\mathscr{V}_{x}}$ and
@@
(a) and (b) are clear.
@@
$=\mathrm{diag}(\langle\xi,\eta\rangle_{\mathcal{E}}(x),\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha(x)),\ldots,\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-1}(x)))(a_{0},a_{1},\ldots,a_{r_{k}-1}).$
@@
$\ (\pi_{k}(f)\tau_{k}(\xi))(x)(a_{0},\dots,a_{r_{k}-1})$
$=$ $\ \pi_{k}(f)(x)(0,\xi(x)\otimes a_{0},\dots,\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ (0,f(\alpha(x))\xi(x)\otimes a_{0},\dots,f(\alpha^{r_{k}-1}(x))\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
@@
$=$ $\ (0,(f\xi)(x)\otimes a_{0},\dots,(f\xi)(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ \tau_{k}(f\xi)(a_{0},\dots,a_{r_{k}-1}).$
@@
$=$ $\ (0,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha(x)),\ldots,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha^{r_{k}-1}(x)))$
@@
On the other hand, with $a_{i}=a_{i}^{\prime}\otimes a_{i}^{(i-1)}\in\mathscr{V}_{\alpha^{i-1}(x)}\otimes\mathscr{V}^{(i-1)}_{x}$, we have, for $11\leq i\leq r_{k}-$ 1,
@@
$=$ $\ (0,\xi(x)\otimes\langle\eta(x),a_{1}\rangle_{\mathscr{V}_{x}},\dots$
$\dots,\xi(\alpha^{r_{k}-2}(x))\otimes\langle\eta(\alpha^{r_{k}-2}(x)),a_{r_{k}-1}^{\prime}\rangle_{\mathscr{V}_{\alpha^{r_{k}-2}(x)}}a_{r_{k}-1}^{(r_{k}-2)}).$
@@
simply because of the fact that $\langle v_{1},v_{2}\rangle_{\mathscr{V}_{y}}v_{3}=\langle v_{1},v_{3}\rangle_{\mathscr{V}_{y}}v_{2}$ for any $y\in X$ and $v_{1},v_{2},v_{3}\in\mathscr{V}_{y}$ which can be seen by considering their inverse images under a chart map $h_{U}$ for $U\in\mathcal{U}$ with $y\in U$. ∎
@@
    which we will denote by $\pi_{k}$ again.
@@
    There is a gauge action $\beta_{z}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k})$ for $z\in\mathbb{T}$.
@@
(a) Let $\varphi_{\mathcal{E}_{Y}}$ be the left action of $C(X)$ on $\mathcal{E}_{Y}$ which is the restriction of the left action of $C(X)$ to $\mathcal{E}$. Since $J_{\mathcal{E}_{Y}}:=\varphi_{\mathcal{E}_{Y}}^{-1}(\mathcal{K}(\mathcal{E}_{Y}))\cap(\ker\varphi_{\mathcal{E}_{Y}})^{\perp}=C_{0}(X\setminus Y)$ (this can be shown easily, we have to show that
@@
where $\psi_{\tau_{k}}:\mathcal{K}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$ is the -homomorphism given by $\psi_{\tau_{k}}(\theta_{\xi,\eta})=\tau_{k}(\xi)\tau_{k}(\eta)^{*}$. But this follows from Lemma 7.2.(e) once we have the following
$\theta_{\xi,\eta}=\varphi_{\mathcal{E}_{Y}}(_{\mathcal{E}}\langle\xi,\eta\rangle),\qquad\xi,\eta\in\mathcal{E}_{Y}$
which actually holds since for any $\zeta\in\mathcal{E}_{Y}$,
Now by the universal property of $\mathcal{O}(\mathcal{E}_{Y})$, there exists a -homomorphism of $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ which we also denote $\pi_{k}$.
(b) For $z\in\mathbb{T}$, let $\beta_{z}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k})$ be the *-homomorphism given by $\beta_{z}=Ad(U_{z})$ on each fiber at $x$ for a unitary map $U_{z}$ which is an endomorphism on each fiber at $x$ by
@@
Thus the representation $(\pi_{k},\tau_{k})$ admits a gauge action for each $z\in\mathbb{T}$. ∎
@@
To see that $\pi$ is injective, we only need to show that it is injective on $C(X)$ by [23, Theorem 6.4] since $\pi(\mathcal{O}(\mathcal{E}_{Y}))$ admits the gauge action
$z\mapsto\beta_{z}:=\beta_{1,z}\oplus\cdots\oplus\beta_{K,z},$
where $z\mapsto\beta_{k,z}$ is the gauge action on $\pi_{k}(\mathcal{O}(\mathcal{E}_{Y}))$ obtained in Lemma 7.3. Let $f\in C(X)$ be a function such that $\pi(f)=\pi_{1}(f)\oplus\cdots\oplus\pi_{K}(f)=0$. Then
@@
=\ = $\pi_{k}(\xi_{1})(x)\pi_{k}(\xi_{2})(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
=\ = $\pi_{k}(\xi_{1})(x)(0,\xi_{2}(x)\otimes a_{0},\dots,\xi_{2}(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
=\ = $(0,0,\xi_{1}(\alpha(x))\otimes\xi_{2}(x)\otimes a_{0},\dots,\xi_{1}(\alpha^{r_{k}-2}(x))\otimes\xi_{2}(\alpha^{r_{k}-3}(x))\otimes a_{r_{k}-3})$
=\ = $(0,0,\psi(\xi_{1}\otimes\xi_{2})(x)\otimes a_{0},\dots,\psi(\xi_{1}\otimes\xi_{2})(\alpha^{r_{k}-3}(x))\otimes a_{r_{k}-3}).$
@@
$=$ $\ (0,\ldots,0,\psi(\xi)(x)\otimes a_{0},\psi(\xi)(\alpha(x))\otimes a_{1},\ldots,\psi(\xi)(\alpha^{r_{k}-1-n}(x))\otimes a_{r_{k}-1-n})$

### v2

@@ 6.0.2. Local trivialization of $\mathscr{M}^{(n)}$
    since $\mathscr{V}^{(i)}_{x}$ and $\mathscr{V}_{\alpha^{i}(x)}^{(m)}$ are all one dimensional. Using the equation (14) we see that
    $\varsigma^{(i)}(x)=s_{m+i,i}(x)v^{(m)}_{\alpha^{i}(\mathbf{U})}(\alpha^{i}(x))$ (16)
    since $v^{(m)}_{\alpha^{i}(\mathbf{U})}(\alpha^{i}(x))=h_{U_{m+i-1}}(\alpha^{m+i-1}(x),1)\otimes\cdots\otimes h_{U_{i}}(\alpha^{i}(x),1)$ as defined in (11).
Note that $\Gamma(\mathscr{M}^{(n)})$ is an algebra with respect to multiplication defined by pointwise composition of endomorphisms at each fibre. For $\varsigma\in\Gamma(\mathscr{M}^{(n)})$ and $x\in\mathbf{U}\cap\mathbf{V}$, we have
$H_{\mathbf{U}}^{(n)}(x,(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))^{*})$ $=\ H_{\mathbf{U}}^{(n)}(x,\widetilde{\varsigma}_{\mathbf{U}}(x)^{*})$
$=\ H_{\mathbf{V}}^{(n)}(x,u_{\mathbf{VU}}\widetilde{\varsigma}_{\mathbf{U}}(x)^{*}u_{\mathbf{VU}}^{*})$
$=\ H_{\mathbf{V}}^{(n)}(x,(u_{\mathbf{VU}}\widetilde{\varsigma}_{\mathbf{U}}(x)u_{\mathbf{VU}})^{*})$
$=\ H_{\mathbf{V}}^{(n)}(x,\widetilde{\varsigma}_{\mathbf{V}}(x)^{*})$
$=\ H_{\mathbf{V}}^{(n)}(x,(H_{\mathbf{V}}^{(n)})^{-1}(\varsigma(x))^{*}),$
hence $\varsigma^{*}$ is well defined by $\varsigma^{*}(x):=H_{\mathbf{U}}^{(n)}(x,(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))^{*})$. The -algebra $\Gamma(\mathscr{M}^{(n)})$ is in fact a $C^{*}$-algebra with respect to the norm
$\|\varsigma\|:=\sup_{x\in X}\|(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}.$
Note that $\|(H_{\mathbf{U}}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}=\|(H_{\mathbf{V}}^{(n)})^{-1}(\varsigma(x))\|_{M_{n}(\mathbb{C})}$, so the norm above does not depend on the choice of the set $\mathbf{U}$.
From now on we fix an infinite compact metric space $X$, a minimal homeomorphism $\alpha:X\to X$, a line bundle $\mathscr{V}$ over $X$, and a closed subset $Y\subset X$ with non-empty interior. Put $\mathcal{E}:=\Gamma(\mathscr{V},\alpha)$ and $\mathcal{E}_{Y}:=C_{0}(X\setminus Y)\mathcal{E}$.
As in Section 5, we denote by $0<r_{1}<\dots<r_{K}$ the finitely many first return times to $Y$ and set
$Y_{k}:=\{y\in Y\mid r_{Y}(y)=r_{k}\}.$
@@ 7. Embedding $\mathcal{O}(\mathcal{E}_{Y})$ into a direct sum of homogeneous algebras
Fix a finite open cover $\mathcal{U}$ of $X$ as in Section 6.0.2. For $k=1,2,\ldots,K$, we define $\mathscr{M}_{k}$ to be the restriction of the endomorphism bundle $\mathscr{M}^{(r_{k})}$ to $\overline{Y_{k}}$, namely
@@
For convenience, we write $H_{\mathbf{U}}^{(k)}:=H_{\mathbf{U}}^{(r_{k})}$, $\mathbf{U}\in\mathcal{U}^{(r_{K})}$, for the rest of the paper.
In this section we show that there exists an embedding $\pi$ of the orbit-breaking algebra $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{1})\oplus\cdots\oplus\Gamma(\mathscr{M}_{K})$, where $\Gamma(\mathscr{M}_{k})$ is the $\mathrm{C}^{*}$-algebra of continuous sections of the endomorphism bundle $\mathscr{M}_{k}$. We begin by constructing a $\mathrm{C}^{*}$-homomorphism $\pi_{k}$ from $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ for every $k=1,\dots,K$. We end the section by proving that $\pi_{1}$ is surjective while it is not true for $\pi_{k}$ when $2\leq k\leq K$ unless $Y_{k}$ is closed.
To construct a *-homomorphism $\pi_{k}$ from the Cuntz–Pimsner algebra $\mathcal{O}(\mathcal{E}_{Y})$ into the $\mathrm{C}^{*}$-algebra $\Gamma(\mathscr{M}_{k})$, we define a covariant representation $(\pi_{k},\tau_{k})$ of $\mathcal{E}_{Y}$ into the $\mathrm{C}^{*}$-algebra $\Gamma(\mathscr{M}_{k})$ which gives rise to the integrated -homomorphism of $\mathcal{O}(\mathcal{E}_{Y})$ into $\Gamma(\mathscr{M}_{k})$ which we will also denote by $\pi_{k}$.
@@
$\pi_{k}(f)(x)(a_{0},\ldots,a_{r_{k}-1})$ $=(f(x)a_{0},f(\alpha(x))a_{1},\ldots,f(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
$\tau_{k}(\xi)(x)(a_{0},\ldots,a_{r_{k}-1})$
$=$ $(0,\xi(x)\otimes a_{0},\xi(\alpha(x))\otimes a_{1},\ldots,\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
for $(a_{0},a_{1},\ldots,a_{r_{k}-1})\in\mathscr{D}^{(r_{k})}_{x}=\mathscr{V}^{(0)}_{x}\oplus\mathscr{V}^{(1)}_{x}\oplus\cdots\oplus\mathscr{V}^{(r_{k}-1)}_{x}$, respectively. Here we mean $\xi(x)\otimes a_{0}:=\xi(x)a_{0}$. By continuity of $f$ and $\xi$, this extends to two maps $\pi_{k}:C(X)\to\Gamma(\mathscr{M}_{k})$ and $\tau_{k}:\mathcal{E}_{Y}\to\Gamma(\mathscr{M}_{k})$.
Let $x\in Y_{k}\cap\mathbf{U}$, $\mathbf{U}=(U_{0},\dots,U_{r_{K}-1})\in\mathcal{U}^{(r_{K})}$, and consider the basis $\{e_{\mathbf{U},0}(x),\dots,e_{\mathbf{U},r_{k}-1}(x)\}$ of $\mathscr{D}^{(r_{k})}_{x}$. If $\xi\in\mathcal{E}_{Y}$, then for each $j=0,\dots,r_{k}-2$, the scalar $\widetilde{\xi}(\alpha^{j}(x))\in\mathbb{C}$ with
$\xi(\alpha^{j}(x))=\widetilde{\xi}(\alpha^{j}(x))h_{U_{j}}(\alpha^{j}(x),1),$ (17)
@@
$\tau_{k}(\xi)(x)e_{\mathbf{U},j}(x)=\widetilde{\xi}(\alpha^{j}(x))e_{\mathbf{U},j+1}(x),$ (18)
@@
$U_{x}(\lambda_{0},\dots,\lambda_{r_{k}-1})=\lambda_{0}e_{\mathbf{U},0}(x)+\cdots+\lambda_{r_{k}-1}e_{\mathbf{U},r_{k}-1}(x),$
we have the matrix representation
@@
This matrix representation will be used in the subsequent discussions along with the fact that the matrix depends on the choice of $\mathbf{U}\in\mathcal{U}^{(r_{K})}$ with $x\in\mathbf{U}$, but the matrix entries differ only by multiplication by scalars of modulus one. For $f\in C(X)$, the corresponding matrix representation is clearly
@@
It is rather obvious that $\pi_{k}$ is in fact a -homomorphism, and we may write $\pi_{k}(f)(x)$ as
@@
where
$\xi(x)^{*}\otimes a_{1}:=\langle\xi(x),a_{1}\rangle_{\mathscr{V}_{x}}$
and
@@ 5. $(e)$
(a) This is clear.
(b) This is also clear.
@@
$=\mathrm{diag}(\langle\xi,\eta\rangle_{\mathcal{E}}(x),\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha(x)),\ldots,\langle\xi,\eta\rangle_{\mathcal{E}}(\alpha^{r_{k}-1}(x)))(a_{0},a_{1},\ldots,a_{r_{k}-1}),$
@@
$\ (\pi_{k}(f)\tau_{k}(\xi))(x)(a_{0},a_{1},\dots,a_{r_{k}-1})$
$=$ $\ \pi_{k}(f)(x)(0,\xi(x)\otimes a_{0},\xi(\alpha(x))\otimes a_{1},\dots,\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ (0,f(\alpha(x))\xi(x)\otimes a_{0},f(\alpha^{2}(x))\xi(\alpha(x))\otimes a_{1},$
$\dots,f(\alpha^{r_{k}-1}(x))\xi(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
@@
$=$ $\ (0,(f\xi)(x)\otimes a_{0},(f\xi)(\alpha(x))\otimes a_{1},\dots,(f\xi)(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=$ $\ \tau_{k}(f\xi)(a_{0},a_{1},\dots,a_{r_{k}-1}).$
@@
$=$ $\ (0,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha(x))a_{1},\ldots,{}_{\mathcal{E}_{Y}}\langle\xi,\eta\rangle(\alpha^{r_{k}-1}(x))a_{r_{k}-1})$
@@
On the other hand, with $a_{i}=a_{i}^{\prime}\otimes a_{i}^{(i-1)}\in\mathscr{V}_{\alpha^{i-1}(x)}\otimes\mathscr{V}^{(i-1)}_{x}$ for $11\leq i\leq r_{k}-$ 1,
@@
$=$ $\ (0,\xi(x)\otimes\langle\eta(x),a_{1}\rangle_{\mathscr{V}_{x}},\dots,\xi(\alpha^{r_{k}-2}(x))\otimes\langle\eta(\alpha^{r_{k}-2}(x)),a_{r_{k}-1}^{\prime}\rangle_{\mathscr{V}_{\alpha^{r_{k}-2}(x)}}a_{r_{k}-1}^{(r_{k}-2)}).$
@@
simply because of the fact that $\langle v_{1},v_{2}\rangle_{\mathscr{V}_{y}}v_{3}=\langle v_{1},v_{3}\rangle_{\mathscr{V}_{y}}v_{2}$ for any $y\in X$ and $v_{1},v_{2},v_{3}\in\mathscr{V}_{y}$ which can be seen by considering their inverse images under a chart map $h_{\mathbf{U}}$ for $\mathbf{U}\in\mathcal{U}^{(1)}$ with $y\in\mathbf{U}$. ∎
@@
    which we also denote by $\pi_{k}$.
@@
    $\Gamma(\mathscr{M}_{k})$ admits a gauge action $\beta:\mathbb{T}\times\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k}),(z,\xi)\mapsto\beta_{z}(\xi)$.
@@
(a) Lemma 7.3 (a)-(d) imply that $(\pi_{k},\tau_{k})$ is a representation of $\mathcal{E}_{Y}$ on $\Gamma(\mathscr{M}_{k})$. It remains to show that the representation is covariant. Let $\varphi_{\mathcal{E}_{Y}}$ be the left action of $C(X)$ on $\mathcal{E}_{Y}$ given by restriction of the structure map $\varphi_{\mathcal{E}}:C(X)\to\mathcal{L}(\mathcal{E})$. Let
$J_{\mathcal{E}_{Y}}:=\varphi_{\mathcal{E}_{Y}}^{-1}(\mathcal{K}(\mathcal{E}_{Y}))\cap(\ker\varphi_{\mathcal{E}_{Y}})^{\perp}.$
We have
for any $\xi,\eta,\zeta\in\mathcal{E}_{Y}$, from which it follows that
$\theta_{\xi,\eta}=\varphi_{\mathcal{E}_{Y}}(_{\mathcal{E}}\langle\xi,\eta\rangle),\qquad\xi,\eta\in\mathcal{E}_{Y}.$
Hence $(\pi_{k},\tau_{k})$ is a covariant representation since Lemma 7.3(e) gives
@@
where $\psi_{\tau_{k}}:\mathcal{K}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$ is the -homomorphism given by $\psi_{\tau_{k}}(\theta_{\xi,\eta})=\tau_{k}(\xi)\tau_{k}(\eta)^{*}$. It follows that $J_{\mathcal{E}_{Y}}=C_{0}(X\setminus Y)$.
Now the existence of the -homomorphism $\pi_{k}:\mathcal{O}(\mathcal{E}_{Y})\to\Gamma(\mathscr{M}_{k})$ follows from the universal property of $\mathcal{O}(\mathcal{E}_{Y})$.
(b) For $z\in\mathbb{T}$, let $\beta_{z}:\Gamma(\mathscr{M}_{k})\to\Gamma(\mathscr{M}_{k})$ be the *-homomorphism given by $\beta_{z}=Ad(U_{z})$ on each fibre $\mathscr{M}^{k}_{x}$ at $x$ for a unitary map $U_{z}$ which is an endomorphism on each fibre $\mathscr{D}^{(r_{k})}_{x}$ at $x$ by
@@
Thus the representation $(\pi_{k},\tau_{k})$ admits a gauge action. ∎
@@
The gauge actions $z\mapsto\beta_{k,z}$ on $\pi_{k}(\mathcal{O}(\mathcal{E}_{Y}))$, $k=1,\dots,K$, obtained in Lemma 7.4, define a gauge action on $\pi(\mathcal{O}(\mathcal{E}_{Y}))$
$z\mapsto\beta_{z}:=\beta_{1,z}\oplus\cdots\oplus\beta_{K,z}.$
Thus to show that $\pi$ is injective, it is enough to show injectivity on $C(X)$ [25, Theorem 6.4].
Let $f\in C(X)$ be a function such that $\pi(f)=\pi_{1}(f)\oplus\cdots\oplus\pi_{K}(f)=0$. Then
@@
$=\pi_{k}(\xi_{1})(x)\pi_{k}(\xi_{2})(x)(a_{0},a_{1},\ldots,a_{r_{k}-1})$
$=\pi_{k}(\xi_{1})(x)(0,\xi_{2}(x)\otimes a_{0},\dots,\xi_{2}(\alpha^{r_{k}-2}(x))\otimes a_{r_{k}-2})$
$=(0,0,\xi_{1}(\alpha(x))\otimes\xi_{2}(x)\otimes a_{0},\dots,\xi_{1}(\alpha^{r_{k}-2}(x))\otimes\xi_{2}(\alpha^{r_{k}-3}(x))\otimes a_{r_{k}-3})$
$=(0,0,\psi(\xi_{1}\otimes\xi_{2})(x)\otimes a_{0},\dots,\psi(\xi_{1}\otimes\xi_{2})(\alpha^{r_{k}-3}(x))\otimes a_{r_{k}-3}).$
@@
$=(0,\ldots,0,\psi(\xi)(x)\otimes a_{0},\psi(\xi)(\alpha(x))\otimes a_{1},\ldots,\psi(\xi)(\alpha^{r_{k}-1-n}(x))\otimes a_{r_{k}-1-n}),$